OpenStudy (anonymous):

Compare the table and equation. X Y 2 10 3 15 4 20 Equation: y = 4.5x Which representation has the greatest slope? (4 points) The table has the greatest slope. The equation has the greatest slope. Their slopes cannot be determined. The table and equation have the same slope.

OpenStudy (anonymous):

The definition of slope is how much y changes divided by how much x changes, or: \[\frac{\Delta y}{\Delta x}\] Try using this formula to calculate the slope for each.

OpenStudy (anonymous):

so... for the table... y:10/2? like that so far right?

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

Ok so 10/2=5, 15/3=5, 20/4=5. so for the table its all five. and the equation is 4.5 and 5 is bigger than 4.5 so the table has the greatest slope

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Right answer, not exactly right on the process. The \(\Delta\) in \(\Delta y\) and \(\Delta x\) stands for the \(\textit{change}\) in \(x\) or \(y\). The change in x in the table is \(4-2 = 2\), because x starts at 2 and moves to 4. The change in y is \(20-10 = 10\). So, \(\frac{\Delta y}{\Delta x}\) becomes \(\frac{10}{2} = 5 \). For the equation, no matter what x is, y is 4.5 times that. So, the change in y is always going to be 4.5 times the change in x, so the slope is 4.5

OpenStudy (anonymous):

@natbelike whyd you block me